# Molecular Dynamics Study of Rayleigh-Plateau Instability at Liquid-Liquid Interfaces

### v1

@@ Abstract
We investigated the Rayleigh-Plateau instability at the interface between two immiscible liquids of equal viscosity using molecular dynamics simulations. We examined two types of initial conditions: one involving a single-mode perturbation at the interface and the other without any imposed perturbation. The growth rate under single-mode perturbation showed strong agreement with classical theory, particularly for large cylinder radii. However, for smaller radii, the growth rate deviated significantly from the theoretical prediction. This discrepancy can be attributed to the effects of thermal fluctuations or pressure differences between the inside and outside of a cylinder. In contrast, with initial conditions free of perturbation, we observed a relationship between breakup time and minimum radius, where the power exponent decreased as the radius decreased. This result suggests that thermal fluctuations have a more substantial impact on systems of smaller dimensions. This study demonstrated that imposing single-mode perturbations effectively suppresses thermal noise and mode coupling, thereby yielding growth rates consistent with those predicted by linear stability analysis.
@@ I Introduction
The particle-based numerical methods, such as dissipative particle dynamics and molecular dynamics (MD), allow us to investigate the behavior of the breakup on a molecular scale. These approaches can also quantify detailed physical properties inaccessible to experiments, enabling direct comparisons with theoretical predictions. The numerical studies have examined the cylindrical geometry near breakup, the size distribution of droplets post-breakup, and the influence of surfactants on destabilizationGopan and Sathian (2014); Tiwari and Abraham (2008); Carnevale et al. (2024). In particular, thermal fluctuations have been found to affect the pre-breakup interface shape, indicating that fluid equations incorporating thermal fluctuation terms become dominant on microscopic scalesMoseler and Landman (2000); Zhao, Sprittles, and Lockerby (2019).
However, the gas and liquid phases are combined after being separately equilibrated in order to investigate interfacial instabilities. This procedure complicates the analysis, as the resulting instabilities are influenced by evaporation.
@@
In the present study, we investigated the instability of the interface between two liquids with identical physical properties by MD simulations. We determined the growth rate to compare with the linear stability theory. We first performed simulations of single-component fluids to prepare the system with the liquid-liquid interface and obtained the equilibrium state. We then generated liquid-liquid interfaces from the resulting single-component liquids by setting different types of atoms inside and outside the cylinder shape. In this way, we achieved conditions in which properties are identical inside and outside the interface, satisfying the conditions required by Stone’s theory. Therefore, we can directly compare the simulation results to the theory. Furthermore, the interface between liquids that do not mix is more stable than the gas-liquid interface and is not affected by anything other than RP instability.
We considered two types of initial conditions: one with a single-mode perturbation imposed on the interface, and the other without any imposed perturbation. With a single-mode perturbation as the initial condition, the growth rate of the mode exhibited good agreement with Stone’s theory in the region of a large cylinder radius. In contrast, in the small-radius regime, the growth rate was significantly larger than the theoretical prediction. To further investigate this, we examined the relationship between breakup time and minimum radius under unperturbed initial conditions. The results showed that the power exponent decreased as the radius decreased, indicating that thermal fluctuations have a stronger influence at smaller scales. Taken together, these findings suggest that imposing a single-mode perturbation can suppress the effects of thermal fluctuations and mode coupling, thereby enabling the mode growth rates to align more closely with predictions from linear stability analysis.
@@ II.1 Linear stability analysis
Stone and Brenner described the linear stability analysis of the RP instability for fluids with equal viscositiesStone and Brenner (1996). When thermal fluctuations are not considered, previous studies have shown that the dimensionless growth rate $\tilde{\omega}(\tilde{k})$ can be described as the dimensionless wavenumber $\tilde{k}=r_{0}k$:
$\tilde{\omega}(\tilde{k})=\frac{\mu r_{0}}{\gamma}\omega(\tilde{k})=(1-\tilde{k}^{2})\left[I_{1}(\tilde{k})K_{1}(\tilde{k})+\frac{\tilde{k}}{2}(I_{1}(\tilde{k})K_{0}(\tilde{k})-I_{0}(\tilde{k})K_{1}(\tilde{k}))\right],$ (1)
@@
Then we determined dimensionless growth rate $\tilde{\omega}(\tilde{k})$ by the interfacial tension $\gamma$ and viscosity $\mu$, which we determined separately.
@@ II.2 Molecular dynamics simulations
where $\sigma$ is the atomic diameter, $\varepsilon$ is the interaction energy, $r$ is the interatomic distance, and $r_{\mathrm{c}}$ is the cutoff distanceLennard-Jones (1931); Weeks, Chandler, and Andersen (1971), respectively. In the following, we adopted the reduced unit such that the $\sigma,\varepsilon$, and $k_{\mathrm{B}}$ are unity. We set the cutoff $r_{\mathrm{c}}=3.0$ for interactions between atoms of the same type to represent attractive interactions. On the other hand, we set $r_{\mathrm{c}}=2^{1/6}\approx 1.12246$ for atoms with different types, which is the Weeks-Chandler-Andersen potential, to represent repulsive interactions. We used the velocity-Verlet algorithm with a time step of 0.005Verlet (1967). The simulations were performed in the NVT ensemble. The size of the simulation box is $(L_{x},L_{y},L_{z})=(100,100,1000)$. The total number of atoms is $8000000$. We defined the axial direction of the cylinder to be the $z$ direction. The temperature was maintained with Langevin thermostatPaquet, Viktor et al. (2015); Schneider and Stoll (1978). We performed the MD simulations using the Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS)Thompson et al. (2022).
@@ II.4 Perturbed configuration
Figure 3: The schematic view of the cylinder with the imposed perturbation. (a) The schematic view of the cylinder at $t=0$ with the radius $r_{0}$, the amplitude $A$ and the wavelength of the perturbation $\lambda$. (b) The schematic view of the cylinder at time $t$. If only the perturbation with wavelength $\lambda$ grows, the amplitude evolves as $A\exp(\omega t)$ over time.
@@ III.1 Perturbed Configuration
The method of determining the growth rate from the breakup time exhibits the strong finite-size effect and could not capture the negative growth rates of stable modes. Therefore, we instead determined the growth rate from the time evolution of the amplitude of the specified wavenumber. Figure 6 shows the time evolution of the amplitude $A(t)$ of the imposed wavenumber $\tilde{k}$. Since the amplitude $A(t)$ grows exponentially, we determine the growth rate $\omega(\tilde{k})$ assuming the relation Eq. (8). Repeating this procedure for various wavenumbers, we obtained the wavenumber dependence of the growth which is shown in Fig. 7.
@@ III.2 Unperturbed Configuration
Assuming the relation $r_{\mathrm{min}}\propto(t_{c}-t)^{\alpha}$, we determined the exponent $\alpha$. We evaluated this exponent for various initial radii. The initial radius dependence of the exponent is shown in Fig. 9. It was found that the value of the exponent increases as the initial radius increases. Since a stronger influence of thermal fluctuations is expected to result in a smaller exponent, this is consistent with the theoretical predictionEggers (1993, 2002); Gopan and Sathian (2014).
@@ IV Discussion and Summary
We performed MD simulations to compare RP instability and linear stability theory. By preparing a liquid-liquid interface as the initial state, we were able to investigate instability without being affected by evaporation, which occurs at a gas-liquid interface. To the best of our knowledge, this is the first direct verification of Stone’s theoretical prediction, which was derived under the condition that the physical properties of the fluid inside and outside the cylinder are identical.
We first observed the time evolutions of the interfaces from initial conditions with disturbances of specified wavenumbers. We considered two observables: the breakup time and the time evolution of the amplitude of the imposed disturbance. The growth rates estimated from the two observables approach those predicted by the classical theory as the radius of the cylinder increases. In particular, the growth rate calculated from the time evolution of the amplitude was almost perfectly consistent with the classical theory at $r_{0}=15$. The fact that thermal fluctuations are practically negligible and consistent with a continuum description in a system as small as 15 atoms in diameter is remarkable.
We also considered the RP instability at the unperturbed initial condition. We observed the time evolution of the minimum radius $r_{\mathrm{min}}$. We assumed the power-law behavior $r_{\mathrm{min}}\propto(t_{c}-t)^{\alpha}$ where $t_{c}$ is the breakup time. The exponent $\alpha$ increased as the radius increased. This strong finite-size effect suggests that the impact of thermal fluctuations is prominent in this system. Nevertheless, the results with the single-mode perturbation, having a relatively small amplitude, agreed well with the classical theory, which neglects the effects of thermal fluctuations. This result indicates that single-mode perturbations effectively suppress thermal fluctuations, allowing instabilities at the microscale to be well described by the classical theory at the macroscale.
A remaining issue is to quantify the influence of thermal fluctuations. While this study clarified that thermal fluctuations cause deviations from classical theory, the extent of that effect remains unclear. By quantifying the effect of thermal fluctuations, the validity of hydrodynamic equations in microscopic environments may be determined. When the radius is small, the Young-Laplace pressure creates a pressure difference between the inside and outside of the cylinder. Since Stone’s theory discusses the instability at the interface of liquids with the same viscosity, the extent to which this pressure difference affects the results needs to be evaluated separately. Evaluating this effect is another important issue for future work.
@@ Acknowledgements.
The authors would like to thank H. Noguchi and H. Nakano for fruitful discussions. This research was supported by JSPS KAKENHI, Grant No. JP21K11923. The computation was partly carried out using the facilities of the Supercomputer Center, Institute for Solid State Physics (ISSP), University of Tokyo.

### v2

@@ Abstract
We investigated the Rayleigh-Plateau instability at the interface between two immiscible liquids of equal viscosity using molecular dynamics simulations. Two types of initial conditions were considered, one with an imposed single-mode perturbation at the interface and the other without any imposed perturbation. Under the single-mode perturbation, the growth rate deviated from the theoretical prediction for small cylinder radii, but progressively approached and agreed with classical macroscopic theory as the radius increased. In contrast, for the unperturbed initial condition, we found a systematic relationship between the breakup time and the minimum radius, in which the power-law exponent increased with increasing radius. These results demonstrate that, even in extremely microscopic systems with cylinder radii on the order of only about fifteen atomic diameters, the growth of the instability can follow macroscopic theoretical predictions when appropriate conditions are imposed, and that the influence of thermal fluctuations on the breakup dynamics becomes increasingly significant as the system radius decreases.
@@ I Introduction
The particle-based numerical methods, such as dissipative particle dynamics and molecular dynamics (MD), allow us to investigate the behavior of the breakup on a molecular scale. These approaches can also quantify detailed physical properties inaccessible to experiments, enabling direct comparisons with theoretical predictions. The numerical studies have examined the cylindrical geometry near breakup, the size distribution of droplets post-breakup, and the influence of surfactants on destabilizationGopan and Sathian (2014); Tiwari and Abraham (2008); Carnevale et al. (2024). In particular, thermal fluctuations have been found to affect the pre-breakup interface shape, indicating that fluid equations incorporating thermal fluctuation terms become dominant on microscopic scalesMoseler and Landman (2000); Zhao, Sprittles, and Lockerby (2019). The correspondence between theoretical predictions and numerical simulations has also been discussed in previous studies. Eggers et al. explained, using a path-integral approach, that thermal fluctuations reduce the breakup exponent and lead to more abrupt rupture dynamicsEggers (2002), while Gopan et al. evaluated the validity of this theory through numerical simulations of gas-liquid systemsGopan and Sathian (2014).
@@
Previous particle-based simulations of the RP instability have mostly focused on gas-liquid interfaces. However, such systems require separate thermal equilibration of the gas and liquid phases before coupling them, and in addition to interfacial instability, the effects of evaporation must also be taken into account. To address this difficulty, we investigated the instability of the interface between two liquids with identical physical properties by MD simulations. We first performed simulations of single-component fluids to prepare the system with the liquid-liquid interface and obtained the equilibrium state. We then generated liquid-liquid interfaces from the resulting single-component liquids by setting different types of atoms inside and outside the cylinder shape. In this way, we achieved conditions in which properties are identical inside and outside the interface, satisfying the conditions required by Stone’s theory. Therefore, we can directly compare the simulation results to the theory. Furthermore, the interface between liquids that do not mix is more stable than the gas-liquid interface and is not affected by anything other than RP instability.
Molecular dynamics enables phenomena to be resolved from the atomic scale and thus allows direct investigation of how microscopic structures influence macroscopic behavior. However, because the system size is inherently small due to the limitation of the computational resources, the results are affected by the discreteness of particles. In the RP instability, thermal fluctuations play an important role, and in addition, the effects of particle discreteness arising from the finite system size cannot be neglected. To isolate the influence of particle discreteness, we imposed initial interfacial perturbations with amplitudes sufficiently larger than those induced by thermal fluctuations, thereby suppressing thermal effects. We found that the growth rate deviates from Stone’s theoretical prediction at small cylinder radii, whereas good agreement with the theory is observed at larger radii. The fact that good agreement with macroscopic theory is obtained even for extremely small systems with cylinder radii on the order of only about fifteen atomic diameters suggests that the influence of particle discreteness on the RP instability is very small.
To quantitatively investigate the influence of thermal fluctuations on the RP instability, we examined the relationship between the breakup time and the minimum radius under unperturbed conditions. It is known that the minimum radius of a liquid column just before breakup is proportional to a power of the remaining time to breakup, and that the corresponding exponent is affected by thermal fluctuations Eggers (1993); Moseler and Landman (2000); Eggers (2002). By analyzing the breakup dynamics of liquid cylinder with various radii, we investigated the radius dependence of this exponent. We found that the value of the breakup exponent increases as the cylinder radius becomes larger. This result is consistent with the theoretical prediction that the breakup exponent approaches $1$ in the absence of thermal fluctuations.
@@ II.1 Linear stability analysis
[…]
@@
[…]
@@ II.2 Molecular dynamics simulations
where $\sigma$ is the atomic diameter, $\varepsilon$ is the interaction energy, $r$ is the interatomic distance, and $r_{\mathrm{c}}$ is the cutoff distanceLennard-Jones (1931); Weeks, Chandler, and Andersen (1971), respectively. In the following, we adopted the reduced unit such that the $\sigma,\varepsilon$, and $k_{\mathrm{B}}$ are unity. We set the cutoff $r_{\mathrm{c}}=3.0$ for interactions between atoms of the same type to represent attractive interactions. On the other hand, we set $r_{\mathrm{c}}=2^{1/6}\approx 1.12246$ for atoms with different types, which is the Weeks-Chandler-Andersen potential, to represent repulsive interactions. We used the velocity-Verlet algorithm with a time step of 0.005Verlet (1967). The simulations were performed in the NVT ensemble. The size of the simulation box is $(L_{x},L_{y},L_{z})=(100,100,1000)$. The total number of atoms is $8000000$. We defined the axial direction of the cylinder to be the $z$ direction. The temperature was maintained with Langevin thermostatPaquet, Viktor et al. (2015); Schneider and Stoll (1978). Throughout this study, simulations were performed at a temperature of $0.8$. The damping parameter of the Langevin thermostat $\tau_{\mathrm{damp}}$ is set to be $50$. We performed the MD simulations using the Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS)Thompson et al. (2022).
@@ II.4 Perturbed configuration
Figure 3: The schematic view of the cylinder with the imposed perturbation. (a) The schematic view of the cylinder at $t=0$ with the radius $r_{0}$, the amplitude AA and the wavelength of the perturbation $\lambda$. (b) The schematic view of the cylinder at time $t$. If only the perturbation with wavelength $\lambda$ grows, the amplitude evolves as $A\exp(\omega t)$ over time.
@@ III.1 Perturbed Configuration
[…]
@@ III.2 Unperturbed Configuration
Assuming the relation $r_{\mathrm{min}}\propto(t_{c}-t)^{\alpha}$, we determined the exponent $\alpha$. We evaluated this exponent for various initial radii. The initial radius dependence of the exponent is shown in Fig. 9. It was found that the value of the exponent increases as the initial radius increases. Since a stronger influence of thermal fluctuations is expected to result in a smaller exponent, this is consistent with the theoretical predictionEggers (1993, 2002); Gopan and Sathian (2014). Note that values lower than the lower bound reported by Eggers et al. ($\alpha\simeq 0.418$) were observed.
@@ IV Discussion and Summary
We performed MD simulations to compare RP instability and linear stability theory. By preparing a liquid-liquid interface as the initial state, we were able to investigate instability without being affected by evaporation, where the gas-liquid system suffers from it. To the best of our knowledge, this is the first direct verification of Stone’s theoretical prediction, which was derived under the condition that the physical properties of the fluid inside and outside the cylinder are identical.
We first observed the time evolutions of the interfaces from initial conditions with disturbances of specified wavenumbers. We considered two observables, the breakup time and the time evolution of the amplitude of the imposed disturbance. The growth rates estimated from the two observables approach those predicted by the classical theory as the radius of the cylinder increases. In particular, the growth rate calculated from the time evolution of the amplitude was almost perfectly consistent with the classical theory at $r_{0}=15$. The fact that thermal fluctuations are practically negligible and consistent with a continuum description in a system as small as $15$ atoms in diameter is remarkable, and this further suggests that particle discreteness does not strongly influence the RP instability.
We also considered the RP instability at the unperturbed initial condition. We observed the time evolution of the minimum radius $r_{\mathrm{min}}$. We assumed the power-law behavior $r_{\mathrm{min}}\propto(t_{c}-t)^{\alpha}$ where $t_{c}$ is the breakup time. The exponent $\alpha$ increased as the radius increased. This strong finite-size effect suggests that the impact of thermal fluctuations is prominent in this system. Nevertheless, the results with the single-mode perturbation, having a relatively small amplitude, agreed well with the classical theory, which neglects the effects of thermal fluctuations.
A remaining issue is to quantify the influence of thermal fluctuations. While this study clarified that thermal fluctuations cause deviations from classical theory, the extent of that effect remains unclear. By quantifying the effect of thermal fluctuations, the validity of hydrodynamic equations in microscopic environments may be determined. When the radius is small, the Young-Laplace pressure creates a pressure difference between the inside and outside of the cylinder. In addition, the breakup exponent obtained in the unperturbed system was found to be smaller than the value reported by Eggers et al. ($\alpha=0.418$). At present, it is not clear whether this discrepancy originates from the use of a liquid-liquid system, and further investigation is required to clarify this issue. Since Stone’s theory discusses the instability at the interface of liquids with the same viscosity, the extent to which this pressure difference affects the results needs to be evaluated separately. Evaluating this effect is another important issue for future work.
@@ Acknowledgements.
The authors would like to thank H. Noguchi and H. Nakano for fruitful discussions. This research was supported by JSPS KAKENHI, Grant No. JP21K11923. The computati on was partly carried out using the facilities of the Supercomputer Center, Institute for Solid State Physics (ISSP), the University of Tokyo (ISSPkyodo-SC-2025 -Cb-0038).
@@ Appendix B Determining Interfacial Tension and Viscosity
Appendix C Effect of the Langevin Thermostat
Figure 12: Relationship between the damping parameter $\tau_{\mathrm{damp}}$ of the Langevin thermostat and the growth dynamics of the instability amplitude. When the relaxation time is too short ($\tau_{\mathrm{damp}}=10$), the breakup tends to be delayed, whereas for sufficiently large relaxation times ($\tau_{\mathrm{damp}}=50,100$), the variation in the breakup time becomes small.
We used the Langevin thermostat for temperature control. In this method, friction and random forces are applied to atoms, and the temperature is regulated by the balance between these two effects. While the target temperature is determined by the ratio of the friction coefficient to the random force, the relaxation time toward the target temperature is controlled by the friction coefficient. The Langevin thermostat involves two additional forces, the frictional force $F_{f}$ and the random force $F_{r}$. The two forces are expressed as
$F_{f}$ $=-\frac{m}{\tau_{\mathrm{damp}}}v,$ (11)
$F_{r}$ $=\hat{R}\sqrt{\frac{k_{B}Tm}{dt\tau_{\mathrm{damp}}}},$
where $m$ is the mass of the atoms, $v$ is the velocity, $k_{B}$ is the Boltzmann constant, $T$ is a designed temperature, $dt$ is the time step, $\tau_{\mathrm{damp}}$ is the damping parameter, and $\hat{R}$ is the white noise, respectivelyBruenger, Brooks, and Karplus (1984). The damping parameter $\tau_{\mathrm{damp}}$ is the inverse of the friction coefficient and determines the relaxation time to the target temperature. When the friction coefficient is large, atomic motion is strongly suppressed, and therefore the temperature control may influence the breakup dynamics near the pinch-off event.
To examine this effect, we performed simulations with different values of the damping parameter of the Langevin thermostat. Figure 12 shows the relationship between the damping parameter and the temporal evolution of the perturbation amplitude. We found that when the relaxation time was too short, i.e., when the temperature control was excessively strong, the breakup dynamics became slower and the breakup time tended to increase. In contrast, when the relaxation time was increased beyond a certain value, corresponding to weaker thermostatting, the variation in breakup time became small. Although changing the relaxation time also affects the effective viscosity and thus the breakup time does not strictly converge to a unique value, we consider that the relaxation times used in our simulations were sufficiently large and that the reported dynamics are not contaminated by significant thermostat-induced artifacts.
Appendix D Relationship Between Temperature and the Breakup Exponent
Figure 13: Relationship between the breakup exponent and the cylinder radius at different temperatures. No significant change in the exponent was observed with respect to variations in the radius.
The stochastic lubrication equation (SLE) proposed by Moseler et al. includes a thermal fluctuation term in the Navier-Stokes equation, and the coefficient of the thermal noise depends on temperature Moseler and Landman (2000). Eggers et al. pointed out that this thermal fluctuation term leads to a reduction in the breakup exponent. In this section, we performed simulations at different temperatures to investigate the effect of temperature on the breakup exponent. The results are shown in Fig. 13. We found that the radius dependence of the breakup exponent showed little change with temperature.
Lowering the temperature too much leads to freezing of the system. To explore a wider temperature range, it would be necessary to investigate higher-temperature regimes; however, at higher temperatures the interface becomes increasingly diffuse, which in turn requires simulations with larger system sizes.